# A systematic search for big dippers in ASAS-SN

### v1

@@ 3.2 Multi-Dip Eclipsing Binaries
J005437644347 (ASASSN-20nq, Figure 3.1, ASAS-SN Transient Page22footnotemark: 2) is on the giant branch and has four distinct dimming events that are identical to each other. The dips all reach a depth of mag and last for about 8 days. We found a period of days using a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982). The system has km/s and is classified by Gaia as a single lined spectroscopic binary with days, km/s, consistent with the value of , and (Gaia Collaboration et al., 2016, 2023a). Bashi et al. (2022) uses archival RVs and the Gaia RV statistics to predict a score for each SB1 orbit. They report a score of for this orbital solution, which is above their recommended minimum of for a good binary solution. For the Gaia orbit, the binary mass function
and the photometric primary must be at least in order for it to be more massive (i.e. more luminous) than the secondary (). J005437644347 sits high on the giant branch, likely implying a high mass. We used the Gaia orbital solution to predict the times of superior, , and inferior, , conjunctions. Figure 3.2 shows the Gaia SB1 orbit and the times of superior and inferior conjunction predicted by the Gaia orbit compared to the ASAS-SN light curve. While the uncertainties on the predicted conjunction times are small ( days), we find that they do not correspond to the observed eclipses in ASAS-SN. If we adjust the Gaia periastron time by days to align the inferior conjunction times with the observed ASAS-SN eclipses, the superior conjunctions lie in the seasonal gaps of the ASAS-SN data.
@@
J005255633515 shows two dimming events (Figure 3.1). The first event began on 11/08/2019 and lasted for 11 days, with a maximum depth of mag. The second event began on 08/08/2020 and lasted for 8 days, with a maximum depth of mag. A dip in the WISE light curve (Figure 3.2e, mag) coincides with the second eclipse. About 986 days before the first event in the -band data, there is the beginning of an eclipse in the -band data on 28/11/2016, with the remainder falling in a seasonal gap. We are inclined to believe these are eclipses, due to the shape of the feature and its detection in multiple cameras. For days, km/s, so the period seems consistent with Gaia radial velocity amplitude of km/s. J005255633515 is on the giant branch and is similar to J181752580749.
@@
J062510075341 had two deep dimming events and the first began on 14/08/2019 and lasted for about 4 days, with a depth of mag. The second event began on 30/10/2024 and lasted for 3 days, with a depth of mag. The minima of the two events are separated by about 1905 days. The target is a red giant and the event is similar to the event seen in J005437644347 above. Gaia reports a single-lined spectroscopic binary (SB1) orbit solution with an orbital period of days, a velocity semi-amplitude of km/s, and an eccentricity of (Gaia Collaboration et al., 2016, 2023a). This target has a binary orbit score of , which is above the minimum recommended for a good orbital solution by Bashi et al. (2022). Note that the period of days, estimated from Eqn 1 and km/s is consistent with the SB1 period, as well as the period seen in the ASAS-SN light curve, if the mass is fairly high. J062510075341 sits on the giant branch above the red clump, but it is hard to determine the masses for giants. We use the Gaia model to predict the times of superior and inferior conjunctions are and . However, because the Gaia RVs were taken years prior to the observed ASAS-SN event, and the Gaia period uncertainty is days, we cannot make a meaningful comparison between the conjunction times predicted by Gaia and the observed eclipse. The next Gaia data release is expected to include epoch radial velocity measurements, and a simultaneous fit to the Gaia RVs and the ASAS-SN light curve will allow the orbital elements to be better constrained.
@@
J124745622756 shows two deep, dimming events. The first event begins on 10/05/2021 and lasts for about 7 days with a maximum depth of mag. The second event begins on 14/05/2024 and also lasts for about 7 days with the same maximum depth. The two dips are separated by about 1096 days. This target is similar to the Gaia single lined spectroscopic binary J005437644347 (Figure 3.1). J124745622756 is in the Hertzsprung Gap, but the star is in the Galactic plane ( deg) and the 3-dimensional dust maps predict a high extinction ( mag), so the CMD position is uncertain. It has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023a), while for a 1096 day period, km/s. The high velocity amplitude implies a high mass or eccentricity.
@@
J160757574540 is a red clump star with three distinct deep, dimming events in the -band, with one proceeding -band. The first dimming event began on 26/08/2018, with a depth of mag, and lasting for about 3 days. The second dimming event began on 20/06/2021, reached a depth of mag, and lasted for about 4 days. The third dimming event began on 04/05/2023, had a depth of mag, and lasted for about 8 days. The fourth dimming event began on 17/04/2024, reached a maximum depth of mag and lasted for about 2 days. The minima of the first and second event are separated by about 1031 days. The minima of the second and third events are separated by about 706 days, while the minima of the primary eclipses (second and fourth events) are separated by about 1030 days. In total there are four dimming events, but the primary and secondary eclipses are not separated by 515 days, indicating that the binary has a non-zero eccentricity. This target has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023a) which suggests binarity. A 1030 day period implies km/s which requires the source to have either a high mass or eccentricity to have the significantly larger . J160757574540 sits on the giant branch above the red clump, but is difficult to infer the masses for giants.
@@
J175912120956 features two deep, dimming events. The first event takes place on 06/05/2021 with a maximum depth of mag and lasts for less than a day. The second event takes place on 12/05/2022 with a maximum depth of mag and, also lasts for less than a day. The minima of the dips are separated by about 371 days. There was also a dimming event in the V-band that began on 12/04/2017 with a maximum depth of mag and also lasting less than a day. The minima of the -band event and the first -band event are separated by about 1485 days. This is 4 times 371, so this is a likely eclipsing binary that has a period of days. Even though the duration of the eclipses is short, we detect the eclipses in multiple cameras. The first -band event has two points from different cameras. The second -band event has six points from two different cameras. This target has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023a). With a period of days, km/s. The large cannot be due to the long-period orbit, but it is possible that this system is instead a hierarchical triple, with the corresponding to an inner binary and the eclipses being due to the more distant companion. However, J175912120956 is on the upper main sequence, and since Gaia RVs can be poorly measured for hot stars, it is also possible that the RV amplitude estimate is biased.
@@ 3.3 Dippers
J070519061219 (ASASSN-24fw, Figure 3.3, JoHantgen et al., 2024b; Forés-Toribio et al., 2025; Zakamska et al., 2025) is an upper main sequence star with an 8 month long, deep ( mag), dimming event that began on 16/09/2024 and ended in early June 2025. This event was the subject of an extensive follow-up campaign and is characterized in detail in Forés-Toribio et al. (2025) and Zakamska et al. (2025). It was caused by dust obscuring the object, and it has a period of 43.8 years (Nair & Denisenko, 2024; Forés-Toribio et al., 2025; Zakamska et al., 2025).
@@
J085816430955 (ASASSN-25bv, JoHantgen et al., 2025a) is an upper main sequence star with two deep dimming events that are separated by about 29 days. The first dip began on 14/02/2025, lasted for roughly 27 days and reached a maximum depth of mag. The shape of this event was not symmetric, and the egress was also captured in TESS Sector 89. After the minimum, there is a slow rise in flux until 06/04/2025, when it began to fade again and plateaued at a depth of mag from 10/04/2025 to 16/04/2025 before fading by an additional 0.3 mag. The event is still ongoing. This star has a RUWE = 5.44, and the Gaia km/s, suggesting it is a part of a binary system (Gaia Collaboration et al., 2016, 2023a). The , suggests a possible period of days. The shape of the light curve is qualitatively similar to massive stars occulted by disks, such as TT Nor (Rowan et al., 2023), but the system would have a much longer period. We looked at the DASCH light curve (Laycock et al., 2010; Grindlay et al., 2012; Tang et al., 2013) to try to constrain the period, but there is too much scatter to search for earlier events.
J114712621037 (ASASSN-21js, Pramono et al., 2024) is an upper main sequence star with an ongoing, large, deep dimming event that began on 20/11/2020. The maximum depth so far mag. Pramono et al. (2024) explained the event as being due to a large dust ring occulting the star and predicted the event will end in January 2027.
@@
J183153284827 (ASASSN-25cd, JoHantgen et al., 2025b) is a red giant star with a deep dimming event that began on 30/03/2025 and is still fading. It had reached a depth of mag as of 11/07/2025. The current minimum was not detected by our pipeline, because the event is ongoing. Broadband photometry of this target before the dimming is consistent with a K-giant (McCollum & Laine, 2025). As discussed above, the km/s implies an orbit long compared to the day span of the Gaia data.
@@
J042214152530 shows several dimming events that are irregularly spaced in time. The deepest event begins on 20/07/2022 and lasts for about 135 days. It reaches a maximum depth of mag. This target is similar to J005437644347 (ASASSN-20nq). J042214152530 was classified as a YSO by ASAS-SN (Jayasinghe et al., 2018), a ”dubious” variable by ATLAS with a period of days (Heinze et al., 2018), and a rotational variable by ZTF (Qiao et al., 2024). There are two archival LAMOST spectra, taken in on 2016-02-05 and 2017-01-28, with similar RVs ( km/s and km/s, respectively). The extinction corrected color, mag and absolute magnitude, , places it just below the giant branch (Figure 2). The SED (Figure 3.1d) has an IR excess in the 2MASS and WISE filters and the target is brighter in the band than in the band. The target has also gotten brighter in the and filters over the past years with mag, followed by a dip ( mag) coinciding with the optical dip (Figure 3.2d). Chandra (Evans et al., 2024) reports an X-ray detection with a luminosity of erg using the Gaia distance, which is typical of YSOs (Forbrich et al., 2011). The km/s, which suggests a period of days.
@@
J080327261620 (ASASSN-24fa, JoHantgen et al., 2024a) shows a long-lasting, deep dimming event that began on 31/01/2021 and reached a maximum depth of mag on 07/03/2023. Since then the flux has been slowly recovering. We do not see any earlier events. J080327261620 is on the upper main sequence and is similar to J114712621037. J080327261620 is reported by Gaia as a short period variable ( days, Gaia Collaboration et al., 2016, 2023a), while ATLAS assigns it a low confidence period of days (Heinze et al., 2018), which is roughly twice as long as the Gaia period. AAVSO identifies it as Cassiopeiae or a Eri type variable with the Gaia period. There is also a mid-IR excess (Figure 3.2a) that begins when the optical light fades and is steadily increasing ( mag). The WISE color also gets redder as the event progresses, suggesting that this event is caused by newly formed dust.
@@
J174328343315 shows several deep, dimming events around 22/09/2017 and they continue in both the -band data and -band data for about 1000 days. These dimming events share some similarities with J184916473251 (ASASSN-21nn, Figure 3.3). This star was classified as a YSO in ASAS-SN (Jayasinghe et al., 2018), a “dubious” variable in ATLAS (Heinze et al., 2018), a suspected variable star in ZTF (Chen et al., 2020), and as an RS CVn variable in Gaia (Gaia Collaboration et al., 2016, 2023a). Its CMD position in the Hertzsprung Gap is atypical for RS CVn stars. Its parallax uncertainty is small ( mas) and the extinction is only mag, suggesting the CMD position is correct. It was observed by LAMOST (Cui et al., 2012, LAMOST-LRS), and the spectrum shows a double-peaked H emission line (Figure 3.3). This, combined with the WISE infrared excess and variability seen in Figures 3.1a and 3.2c, suggest that there is an accretion disk obscuring the star and producing irregular variability (Carroll et al., 1991). In WISE, both the W1 and W2 light curves fade by mag following the most recent event. The source has an km/s, implying a period of days.
@@
J202402383938 has a long-lasting, deep, dimming event. The light curve began fading on 14/04/2018 and reached a maximum depth ( mag) on 14/07/2020. The flux has been slowly recovering since then. The event is asymmetric and includes a number of sharp features which were separately identified by our search. The first sharp feature appears to be a smaller dipping event beginning on 16/10/2019 and reached a maximum depth ( mag) on 22/10/2019. There was an earlier deep, dipping event in the ASAS-SN -band data (top panel of Figure 3.3). The minima of the two events are separated by about 1,470 days. J202402383938 is classified in ATLAS as a long-period variable with days (Heinze et al., 2018). J202402383938 is in the Hertzsprung Gap. The spectral energy distribution shows an infrared excess in W1 – W4 (Figure 3.1b) and the WISE fluxes also drop by mag ( mag) during the event (Figure 3.2b). This suggests the event is an eclipse by a circumstellar disk composed of large particles. J202402383938 was classified as a Brackett-emission line star in APOGEE (Campbell et al., 2023). One of the three APOGEE spectra is shown in the middle panel of Figure 3.3. It has many broad, double-peaked Brackett emission lines. The first two spectra were taken as the star dimmed and the third spectrum was taken as it brightened (top panel of Figure 3.3). The bottom panel zooms in on the Brackett transitions of the spectra corresponding to , , and . The first two spectra show broad, double-peaked emission lines consistent with the infrared spectra of Be stars (Huang, 1972). In the third spectra, the emission lines are replaced by absorption lines, similar to those seen in the spectra of shell stars (Rivinius et al., 2006). The emission lines of the first two spectra match that of a disk seen edge-on (Waters & Marlborough, 1992). The emission lines likely come from the inner disk, the magnetosphere of the star and the disk wind region of the star (Tambovtseva et al., 2016, 2017).

### v2

@@ 3.2 Multi-Dip Eclipsing Binaries
J005437644347 (ASASSN-20nq, Figure 3.1, ASAS-SN Transient Page22footnotemark: 2) is on the giant branch and has four distinct dimming events that are identical to each other. The light curve of J005437644347 features an additional vertical line that was triggered by a noisy point in the light curve and is considered a false positive. The dips all reach a depth of mag and last for about 8 days. We found a period of days using a Lomb-Scargle periodogram (Lomb, 1976; Scargle, 1982). The system has km/s and is classified by Gaia as a single lined spectroscopic binary with days, km/s, consistent with the value of , and (Gaia Collaboration et al., 2016, 2023b). Bashi et al. (2022) uses archival RVs and the Gaia RV statistics to predict a score for each SB1 orbit. They report a score of for this orbital solution, which is above their recommended minimum of for a good binary solution. For the Gaia orbit, the binary mass function is
and the photometric primary must be at least in order for it to be more massive (i.e. more luminous) than the secondary (). J005437644347 sits high on the giant branch, likely implying a high mass. We used the Gaia orbital solution to predict the times of superior, , and inferior, , conjunctions. Figure 3.2 shows the Gaia SB1 orbit and the times of superior and inferior conjunction predicted by the Gaia orbit compared to the ASAS-SN light curve. While the uncertainties on the predicted conjunction times are small ( days), we find that they do not correspond to the observed eclipses in ASAS-SN. If we adjust the Gaia periastron time by days to align the inferior conjunction times with the observed ASAS-SN eclipses, the superior conjunctions lie in the seasonal gaps of the ASAS-SN data. The TESS light curve shows a clear eclipse in Sector 58. This corresponds to the third eclipse in the ASAS-SN light curve. This event reaches its lowest point on 15/11/2022. The light curve shows a flat bottom for the eclipse, implying the eclipse is a total eclipse. The eclipse duration of 11 days is consistent with the events in the ASAS-SN data.
@@
J005255633515 shows two dimming events (Figure 3.1). The first event began on 11/08/2019 and lasted for 11 days, with a maximum depth of mag. The second event began on 08/08/2020 and lasted for 8 days, with a maximum depth of mag. A dip in the WISE light curve (Figure 3.2e, mag) coincides with the second eclipse. About 986 days before the first event in the -band data, there is the beginning of an eclipse in the -band data on 28/11/2016, with the remainder falling in a seasonal gap. We are inclined to believe these are eclipses, due to the symmetric shape of the events and its detection in multiple cameras. For days, km/s, so the period seems consistent with the Gaia radial velocity amplitude of km/s. J005255633515 is on the giant branch and is similar to J181752580749.
@@
J062510075341 had two deep dimming events and the first began on 14/08/2019 and lasted for about 4 days, with a depth of mag. The second event began on 30/10/2024 and lasted for 3 days, with a depth of mag. The minima of the two events are separated by about 1905 days. The target is a red giant and the event is similar to the event seen in J005437644347 above. Gaia reports a single-lined spectroscopic binary (SB1) orbit solution with an orbital period of days, a velocity semi-amplitude of km/s, and an eccentricity of (Gaia Collaboration et al., 2016, 2023b). This target has a binary orbit score of , which is above the minimum recommended for a good orbital solution by Bashi et al. (2022). Note that the period of days, estimated from Eqn 1 and km/s is consistent with the SB1 period, as well as the period seen in the ASAS-SN light curve, if the mass is fairly high. J062510075341 sits on the giant branch above the red clump, but it is hard to determine the masses of giants. We use the Gaia model to predict that the expected Julian Dates of the superior and inferior conjunctions are and . However, because the Gaia RVs were taken years prior to the observed ASAS-SN event, and the Gaia period uncertainty is days, we cannot make a meaningful comparison between the conjunction times predicted by the Gaia orbit and the observed eclipse. The next Gaia data release is expected to include epoch radial velocity measurements, and a simultaneous fit to the Gaia RVs and the ASAS-SN light curve will allow the orbital elements to be better constrained.
@@
J124745622756 shows two deep, dimming events. The first event begins on 10/05/2021 and lasts for about 7 days with a maximum depth of mag. The second event begins on 14/05/2024 and also lasts for about 7 days with the same maximum depth. The two dips are separated by about 1096 days. This target is similar to the Gaia single lined spectroscopic binary J005437644347 (Figure 3.1). J124745622756 is in the Hertzsprung Gap, but the star is in the Galactic plane ( deg) and the 3-dimensional dust maps predict a high extinction ( mag), so the CMD position is uncertain. It has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023b), while for a 1096 day period, km/s. The high velocity amplitude implies a high mass or eccentricity. The first eclipse egress is seen in the TESS sector 38 light curve at the expected time.
@@
J160757574540 is a red clump star with three distinct deep, dimming events in the -band, with one proceeding -band. The first dimming event began on 26/08/2018, with a depth of mag, and lasting for about 3 days. The second dimming event began on 20/06/2021, reached a depth of mag, and lasted for about 4 days. The third dimming event began on 04/05/2023, had a depth of mag, and lasted for about 8 days. The fourth dimming event began on 17/04/2024, reached a maximum depth of mag and lasted for about 2 days. The minima of the first and second event are separated by about 1031 days. The minima of the second and third events are separated by about 706 days, while the minima of the primary eclipses (second and fourth events) are separated by about 1030 days. In total there are four dimming events, but the primary and secondary eclipses are not separated by 515 days, indicating that the binary has a non-zero eccentricity. This target has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023b) which suggests binarity. A 1030 day period implies km/s which requires the source to have either a high mass or eccentricity to have the significantly larger . J160757574540 sits on the giant branch above the red clump, but is difficult to infer the masses for giants. The second and third eclipses were partially observed by TESS in sectors 39 and 65, respectively.
@@
J175912120956 features four deep, dimming events. The first identified event is a false positive due to noisy data. The first real -band event takes place on 06/05/2021 with a maximum depth of mag and lasts for less than a day. The second event takes place on 12/05/2022 with a maximum depth of mag and, also lasts for less than a day. The minima of the dips are separated by about 371 days. The third -band takes place on 28/05/2024 with a maximum depth of mag, also lasts for less than a day, and has a separation of 747 days from the previous event. The fourth -band takes place on 30/05/2025 with a maximum depth of mag, also lasts for less than a day, and has a separation of 367 days from the previous -band event. There was also a dimming event in the -band that began on 12/04/2017 with a maximum depth of mag and also lasting less than a day. There appears to be a combined -band/-band event after this -band event, however, it is not real and consists of several points at the mean magnitude of the light curve in between the apparent points of the event. The nonphysical shape of this event leads us to believe it is a false positive. The minima of the -band event and the first -band event are separated by about 1485 days. This is 4 times 371, so this is a likely eclipsing binary that has a period of days, consistent with the separations of the -band dips. Even though the duration of the eclipses is short, we detect the eclipses in multiple cameras. The first -band event has two points from different cameras. The second -band event has six points from two different cameras. The TESS light curve shows the dip in sector 92 corresponding to the fourth -band dip in the ASAS-SN light curve. Its short duration of days is consistent with the duration of ASAS-SN results. This target has a radial velocity amplitude of km/s in Gaia (Gaia Collaboration et al., 2016, 2023b). With a period of days, km/s. The large cannot be due to the long-period orbit, but it is possible that this system is instead a hierarchical triple, with the corresponding to an inner binary and the eclipses being due to the more distant companion. However, J175912120956 is on the upper main sequence, and since Gaia RVs can be poorly measured for hot stars, it is also possible that the RV amplitude estimate is biased.
@@ 3.3 Dippers
J070519061219 (ASASSN-24fw, Figure 3.3, JoHantgen et al., 2024a) is an upper main sequence star with an 8 month long, deep ( mag), dimming event that began on 16/09/2024 and ended in early June 2025. This event was the subject of an extensive follow-up campaign and is characterized in detail in Forés-Toribio et al. (2025) and Zakamska et al. (2025). It was caused by dust obscuring the object, and it has a period of 43.8 years (Nair and Denisenko, 2024; Forés-Toribio et al., 2025; Zakamska et al., 2025).
@@
J085816430955 (ASASSN-25bv, JoHantgen et al., 2025a) is an upper main sequence star with two deep dimming events that are separated by about 29 days. The first dip began on 14/02/2025, lasted for roughly 27 days and reached a maximum depth of mag. The shape of this event was not symmetric, and the egress was also captured in TESS Sector 89. After the minimum, there is a slow rise in flux until 06/04/2025, when it began to fade again and plateaued at a depth of mag from 10/04/2025 to 16/04/2025 before fading by an additional 0.3 mag. The event is still ongoing. This star has a RUWE = 5.44, and the Gaia km/s, suggesting it is a part of a binary system (Gaia Collaboration et al., 2016, 2023b). The , suggests a possible period of days. The shape of the light curve is qualitatively similar to massive stars occulted by disks, such as TT Nor (Rowan et al., 2023), but the system would have a much longer period. We looked at the DASCH light curve (Laycock et al., 2010; Grindlay et al., 2012; Tang et al., 2013) to try to constrain the period, but there is too much scatter in the data to search for earlier events.
J114712621037 (ASASSN-21js, Pramono et al., 2024) is an upper main sequence star with an ongoing, large, deep dimming event that began on 20/11/2020. The maximum depth so far is mag. Pramono et al. (2024) explained the event as being due to a large dust ring occulting the star and predicted the event will end in January 2027.
@@
J183153284827 (ASASSN-25cd, JoHantgen et al., 2025b) is a red giant star with a deep dimming event that began on 30/03/2025 and is still fading. It reached a depth of mag as of 15 January 2026. Broadband photometry of this target before the dimming is consistent with a K-giant (McCollum and Laine, 2025). As discussed above, the km/s implies an orbit long compared to the day span of the Gaia data.
@@
J042214152530 shows several dimming events that are irregularly spaced in time. The deepest event begins on 20/07/2022 and lasts for about 135 days. It reaches a maximum depth of mag. This target is similar to J005437644347 (ASASSN-20nq). J042214152530 was classified as a YSO by ASAS-SN (Jayasinghe et al., 2018), a ”dubious” variable by ATLAS with a period of days (Heinze et al., 2018), and a rotational variable by ZTF (Qiao et al., 2024). There are two archival LAMOST spectra, taken in on 2016-02-05 and 2017-01-28, with similar RVs ( km/s and km/s, respectively). The extinction corrected color, mag and absolute magnitude, , places it just below the giant branch (Figure 2). The SED (Figure 3.1d) has an IR excess in the 2MASS and WISE filters and the target is brighter in the band than in the band. The target has also gotten brighter in the and filters over the past years with mag, followed by a dip ( mag) coinciding with the optical dip (Figure 3.2d). Chandra (Evans et al., 2024) reports an X-ray detection with a luminosity of erg using the Gaia distance, which is typical of YSOs (Forbrich et al., 2011). The km/s suggests a period of days.
@@
J080327261620 (ASASSN-24fa, JoHantgen et al., 2024b) shows a long-lasting, deep dimming event that began on 31/01/2021 and reached a maximum depth of mag on 07/03/2023. Since then the flux has been slowly recovering. We do not see any earlier events. J080327261620 is on the upper main sequence and is similar to J114712621037. J080327261620 is reported by Gaia as a short period variable ( days, Gaia Collaboration et al., 2016, 2023b), while ATLAS assigns it a low confidence period of days (Heinze et al., 2018), which is roughly twice as long as the Gaia period. AAVSO identifies it as Cassiopeiae or a Eri type variable with the Gaia period. There is also a mid-IR excess (Figure 3.2a) that begins when the optical light fades and is steadily increasing ( mag). The WISE color also gets redder as the event progresses, suggesting that this event is caused by newly formed dust.
@@
J174328343315 shows several deep, dimming events around 22/09/2017 and they continue in both the -band data and -band data for about 1000 days. These dimming events share some similarities with J184916473251 (ASASSN-21nn, Figure 3.3). This star was classified as a YSO in ASAS-SN (Jayasinghe et al., 2018), a “dubious” variable in ATLAS (Heinze et al., 2018), a suspected variable star in ZTF (Chen et al., 2020), and as an RS CVn variable in Gaia (Gaia Collaboration et al., 2016, 2023b). Its CMD position in the Hertzsprung Gap is atypical for RS CVn stars. Since Hertzsprung gap stars are rare, its CMD location may imply that it is an unresolved binary with a red primary and a blue secondary. Its parallax uncertainty is small ( mas) and the extinction is only mag, suggesting the CMD position is correct. It was observed by LAMOST (Cui et al., 2012, LAMOST-LRS), and the spectrum shows a double-peaked H emission line (Figure 3.3). This, combined with the WISE infrared excess and variability seen in Figures 3.1a and 3.2c, suggest that there is an accretion disk obscuring the star and producing irregular variability (Carroll et al., 1991). In WISE, both the W1 and W2 light curves fade by mag following the most recent event. The source has an km/s, implying a period of days.
@@
J202402383938 has a long-lasting, deep, dimming event. The light curve began fading on 14/04/2018 and reached a maximum depth ( mag) on 14/07/2020. The flux has been slowly recovering since then. The event is asymmetric and includes a number of sharp features which were separately identified by our search. The first sharp feature appears to be a smaller dipping event beginning on 16/10/2019 and reached a maximum depth ( mag) on 22/10/2019. There was an earlier deep, dipping event in the ASAS-SN -band data (top panel of Figure 3.3). The minima of the two events are separated by about 1,470 days. J202402383938 is classified in ATLAS as a long-period variable with days (Heinze et al., 2018). J202402383938 is in the Hertzsprung Gap. The spectral energy distribution shows an infrared excess in W1 – W4 (Figure 3.1b) and the WISE fluxes also drop by mag ( mag) during the event (Figure 3.2b). This suggests the event is an eclipse by a circumstellar disk composed of large particles. J202402383938 was classified as a Brackett-emission line star in APOGEE (Campbell et al., 2023). One of the three APOGEE spectra is shown in the middle panel of Figure 3.3. It has many broad, double-peaked Brackett emission lines. The first two spectra were taken as the star dimmed and the third spectrum was taken as it brightened (top panel of Figure 3.3). The bottom panel zooms in on the Brackett transitions of the spectra corresponding to , , and . The first two spectra show broad, double-peaked emission lines consistent with the infrared spectra of Be stars (Huang, 1972). In the third spectra, the emission lines are replaced by absorption lines, similar to those seen in the spectra of shell stars (Rivinius et al., 2006). The emission lines of the first two spectra match that of a disk seen edge-on (Waters and Marlborough, 1992). The emission lines likely come from the inner disk, the magnetosphere of the star and the disk wind region of the star (Tambovtseva et al., 2016, 2017).